KIM SPENT TWO FIFTHS OF HER SALARY ON A SHIRT AND A HAT. THE SHIRT COST THREE TIMES AS MUCH AS THE HAT. IF SHE HAS $36.00 LEFT, HOW MUCH DID THE HAT COST?
Answer by stanbon(48535)

KIM SPENT TWO FIFTHS OF HER SALARY ON A SHIRT AND A HAT. THE SHIRT COST THREE TIMES AS MUCH AS THE HAT. IF SHE HAS $36.00 LEFT, HOW MUCH DID THE HAT COST?
----
Her salary = s + h + 36
----------------------------
s = 3h
s+h = (2/5)[s+h+36]
--------------------------
Substitute to solve for "h":
3h + h = (2/5)[3h + h + 36]
4h = (2/5)[4h + 36]
20h = 2[4h+36]
10h = 4h + 36
6h = 36
h = $6 (the hat cost)
============================
Cheers,
Stan H.